# Group-Aware Partial Model Merging for Children’s Automatic Speech Recognition

### v1

@@ V-A Group-aware model merging
To evaluate the effectiveness of our proposed approach, we first examine whether clustering the training dataset followed by model merging improves performance in children’s ASR, and which utterance embedding strategy yield the best results. As shown in Table II, the baseline system achieves a Word Error Rate (WER) of 14.05%. Fine-tuning the full model on the entire dataset significantly reduces the WER to 9.95%. This result confirms that SFT effectively mitigates the adult-speech bias inherent in the original pre-trained Whisper model.
Next, we partition $\mathcal{D}$ by group using several clustering strategies and apply model merging. For each configuration of $\mathcal{M}$, the different selected fine-tuned models used for merging are listed in the leftmost columns of Table II. Among all evaluated methods, grouping using speaker-embedding (Spk-emb) achieved the overall strongest performance, with an average WER of 9.59% and the lowest WER configuration with a score of 9.36%. The Low-level Speech Metrics (LSM) embeddings clustering also performs well but with less consistent results, reaching an average WER of 9.74% and a best configuration at 9.41%. Random clustering achieves 9.84% on average and 9.41% at best. Meanwhile, the Zero-Shot (ZS) WER clustering performs the least effectively with an average of 10.03% WER and best at 9.86%, showing no improvement over the baseline.
In the configurations where all the fine-tuned models trained on all the different data clusters are merged together (final row of Table II), or $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$, Spk-emb and Random embedding yield slightly higher WERs (9.65%) compared with LSM (9.63%). Nevertheless, all of them outperform the fine-tuning baseline, indicating that model merging remains an effective strategy.
Overall, these findings demonstrate that model merging enhances system performance relative to fine-tuning on the entire dataset and in particular, group-aware with speaker-embedding achieves the strongest and most consistent improvements.
V-B Influence of the number of groups
@@
TABLE III: Influence of the number of clusters combination of all fine-tuned model ($\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$) and the best combination of $\mathcal{M}$ and results in WER (%).
@@
We next evaluate model merging with varying numbers of clusters. Table III presents both the best-performing configurations of $\mathcal{M}$ and the results obtained by merging all fine-tuned group-based models ($\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$) using the Spk-emb clustering method. The single-cluster setting corresponds to training on the entire dataset. For the best configurations, performance improves consistently as the number of clusters increases, with the lowest WER achieved using four clusters with a score of 9.33%. When all group-based models are merged, the combined system also delivers strong results, remaining competitive with the best configuration performance, achieving 9.42% with two clusters. These findings suggest that increasing cluster granularity allows the model to more effectively capture speaker-specific acoustic characteristics, leading to improved ASR performance when selected the right models to merge.
V-C Partial model merging
TABLE IV: WER (%) results of GRAPAM with partial fine-tuning. Grey indicates performance better than both full and partial fine-tuning on $\mathcal{D}$; Diagonal lines pattern indicates improvement over full fine-tuning but not over partial fine-tuning; The best-performing combination is shown in bold.
$\mathcal{M}$ Fine-tuning method
$\theta_{all}$ $\theta_{1}$ $\theta_{2}$ $\theta_{3}$ Full SFT PFT FFN PFT ATTN
x 9.95
Partial 9.48 9.46
x 10.34 10.03 10.00
x 9.79 9.68 9.68
x 10.36 9.58 9.71
x x 9.48 9.52 9.61
x x 9.77 9.69 9.68
x x 9.41 9.44 9.56
x x x 9.68 9.45 9.60
x x 9.72 9.66 9.66
x x 9.41 9.32 9.62
x x 9.76 9.66 9.66
x x x 9.36 9.33 9.68
x x x 9.68 9.32 9.66
x x x x 9.65 9.31 9.38
Partial x x x 9.36 9.51
Average 9.59 9.47 9.60
Table IV reports the WER performance of GRAPAM under three fine-tuning strategies: full model fine-tuning (Full SFT), partial fine-tuning of feed-forward modules (PFT FFN), and partial fine-tuning of attention modules (PFT ATTN). When applied to the entire dataset $\mathcal{D}$, full fine-tuning yields a WER of 9.95%. In contrast, partial fine-tuning achieves superior performance, with WERs of 9.48% (FFN) and 9.46% (ATTN), despite involving fewer trainable parameters.
Applying group-aware model merging further improves performance across all configurations, achieving the lowest WERs of 9.31% with PFT FFN and 9.38% with PFT ATTN. On average, PFT FFN delivers the strongest results (9.47%), outperforming both Full SFT (9.59%) and PFT ATTN (9.60%). Furthermore, we examine whether merging fully fine-tuned models or partially fine-tuned models, both trained on the entire data set, yields greater benefits within the GRAPAM framework. As shown in the final two rows of Table IV, full fine-tuning yields slightly better performance, achieving a WER of 9.31%, compared to 9.36% for PFT–FFN and 9.51% for PFT–ATTN. Interestingly, the best-performing PFT configurations correspond to merging all fine-tuned models, i.e., $\mathcal{M}=\{\theta_{\text{all}},\theta_{1},\theta_{2},\theta_{3}\}$.
V-D Integration of Clustering and Fine-Tuning Strategies
In this section, we explore the integration of the previously discussed clustering methods and fine-tuning strategies, assessing whether performance can be enhanced by merging models across different strategies or by combining them sequentially across training stages.
V-D1 Heterogeneous merging
Table V presents the WER results of GRAPAM for all combination of PFT and utterance embedding when merging all models from all clusters, where $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$. The table also includes the heterogeneous merging scenarios (MERGE ALL line and column), where models fine-tuned using different embedding types and fine-tuning strategies are combined together. The results indicate that PFT–FFN consistently yields the strongest performance with a score of 9.36% WER. Merging over utterance embedding strategy also performs competitively, recording a WER of 9.38% for both speaker embeddings and 9.40% for LSM and 9.39% for random clustering. Importantly, merging across the different utterance embeddings or fine-tuning strategies does not significantly improves or degrades performance, suggesting that the merging process effectively preserves model capabilities.
Notably, the MERGE ALL configuration, which merges all models across clustering types and fine-tuning strategies, achieves a WER of 9.32%, matching the best individual configurations and demonstrating the robustness of the GRAPAM framework under heterogeneous integration.
@@ V-D1 Heterogeneous merging
TABLE V: WER (%) results of different combination of PFT and utterance embedding as well as the heterogeneous merging of them. The best-performing combination is shown in bold.
Full SFT PFT FFN PFT Attn MERGE ALL
Spk embedding 9.65 9.31 9.38 9.38
@@
MERGE ALL 9.60 9.36 9.51 9.32
V-D2 Iterative Group-Aware Partial Merging
TABLE VI: WER (%) for different embedding methods (Spk embedding, LSM, Random) across PFT (FFN, ATTN, FULL) and iterations. Selected models as a source for the next turn are presented in bold.
PFT Utterance embedding TURN 1 TURN 2 TURN 3
Baseline - 9.48 10.14 10.28
@@ V-D2 Iterative Group-Aware Partial Merging
Table VI presents the WER results obtained from successive iterations of GRAPAM, where, at each step, the best-performing configuration from the previous iteration is used as the pre-trained model for subsequent merging. In this experiment we used $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$.
First, we observe that when the full model is trained for additional epochs on the entire dataset, signs of overfitting emerge, as indicated by an increase in WER from 9.48% to 10.28%.
In contrast, the Iterative GRAPAM results demonstrate that initialising with speaker-embedding clustering and partial fine-tuning of the FFN yields the lowest WER of 9.28%, achieved in iteration 2 using LSM-based clustering and full fine-tuning. We hypothesise that this improvement arises from the model’s exposure to heterogeneous information across different cluster types at each iteration, enabling the progressive development of richer and more generalisable representations.
However, beyond iteration 2, no further improvements are observed. This plateau suggests that the model has saturated the informative capacity of the available cluster partitions and may be beginning to overfit.
These findings confirm that iterative group-aware partial model merging effectively enhances ASR performance.
VI Conclusion and future work
In this work, we presented GRAPAM, a parameter-efficient framework for adapting adult-pretrained ASR models to children’s speech through group-aware partial model merging. By combining unsupervised clustering, partial fine-tuning, and parameter-level interpolation, GRAPAM achieves a WER reduction from 9.95% to 9.31% on the MyST corpus. This new fine-tuning strategy for children’s ASR achieves comparable or better performance without needing multi-task training or additional data.
Future work can extend GRAPAM in several directions. First, adaptive weighting, where interpolation coefficients are optimised on held-out validation data rather than set uniformly is a promising direction. Second, more advanced clustering techniques can be considered, potentially incorporating self-supervised speech representations to better capture age and speaker-related variability. Finally, GRAPAM could be extended to other domains like pathological or multilingual speech recognition.

### v2

@@ 5.1 Group-aware model merging
To assess our approach, we first test whether clustering $\mathcal{D}$ followed by model merging improves children’s ASR, and which utterance representation is most effective. As shown in Table 2, the baseline yields 14.05% WER, while full-model fine-tuning on $\mathcal{D}$ reduces WER to 9.95%, confirming that SFT mitigates the adult-speech bias of the pretrained Whisper model.
We then cluster $\mathcal{D}$ using several embedding strategies and merge the resulting domain models according to $\mathcal{M}$. Speaker-embedding (Spk-emb) clustering performs best overall, achieving 9.59% average WER and a best configuration of 9.36%. Low-level speech metrics (LSM) is competitive but less consistent (9.74% average; 9.41% best), while Random clustering is weaker (9.84% average; 9.41% best). Zero-shot WER clustering performs worst (10.03% average; 9.86% best), failing to improve over the fine-tuning baseline.
When merging all cluster-specific models ($\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$), Spk-emb and Random reach 9.65% WER and LSM 9.63%; all remain better than the 9.95% fine-tuning baseline, supporting model merging as an effective consolidation strategy.
5.2 Partial model merging
In Table 2 we compare three fine-tuning strategies: full-model fine-tuning (Full SFT), partial fine-tuning of FFN blocks (PFT FFN), and partial fine-tuning of attention blocks (PFT ATTN). Fine-tuning on the full dataset $\mathcal{D}$ yields 9.95% WER with Full SFT, while partial fine-tuning performs better at 9.48% (FFN) and 9.46% (ATTN), despite fewer trainable parameters.
Group-aware model merging further reduces WER across configurations, reaching best scores of 9.31% (PFT FFN) and 9.38% (PFT ATTN). On average, PFT FFN is strongest (9.47%), outperforming Full SFT (9.59%) and PFT ATTN (9.60%). When merging models trained on the full dataset, fully fine-tuned checkpoints slightly outperform partially fine-tuned ones (9.31% vs. 9.36% for PFT–FFN and 9.51% for PFT–ATTN). The best PFT results arise when merging all cluster models, $\mathcal{M}=\{\theta_{\text{all}},\theta_{1},\theta_{2},\theta_{3}\}$.
5.3 Influence of the number of groups
@@
Table 3: Influence of the number of clusters combination of all fine-tuned model ($\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{k}\}$) and the best combination of $\mathcal{M}$ and results in WER (%).
@@
Table 3 analyses the effect of the number of clusters (Spk-emb) on GRAPAM, reporting both (i) merging all group models $\mathcal{M}=\{\theta_{all},\theta_{1},\dots,\theta_{K}\}$ and (ii) the best-performing subset of $\mathcal{M}$. The K=1 setting (no clustering) matches full-data fine-tuning at 9.95% WER. For the best subsets, increasing K improves performance, reaching the lowest WER with four clusters (9.33%). Merging all group models remains competitive, achieving its best result with two clusters (9.42%). Overall, finer cluster granularity appears to better capture speaker-related variability, yielding gains when the merged model set is appropriately selected.
5.4 Heterogeneous merging
Table 4 reports GRAPAM results when merging all cluster models, $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$, across PFT variants and utterance-embedding choices. It also includes heterogeneous settings (ALL row/column), where models trained with different clustering representations and fine-tuning strategies are merged. Overall, PFT–FFN yields the lowest WER (9.36%). Merging across embedding types remains competitive, with WERs of 9.38% (speaker embeddings), 9.40% (LSM), and 9.39% (random). Importantly, combining models across embeddings or fine-tuning strategies produces only minor changes, indicating that merging largely preserves the underlying capabilities of the constituent models. The full heterogeneous merge (ALL) achieves 9.32% WER, matching the best individual configurations and demonstrating robustness to heterogeneous merge.
@@ 5.4 Heterogeneous merging
Table 4: WER (%) results of different combination of PFT and utterance embedding as well as the heterogeneous merging of them. The best-performing combination is shown in bold.
Full SFT PFT FFN PFT Attn ALL
Spk-emb 9.65 9.31 9.38 9.38
@@
ALL 9.60 9.36 9.51 9.32
5.5 Iterative Group-Aware Partial Merging
Table 5: WER (%) for different embedding methods (Spk embedding, LSM, Random) across PFT (FFN, ATTN, FULL) and iterations. Selected models as a source for the next turn are presented in bold.
PFT Utt embedding Iter 1 Iter 2 Iter 3
Partial - 9.48 10.14 10.28
@@ 5.5 Iterative Group-Aware Partial Merging
Table 5 reports WER for successive GRAPAM iterations, where the best configuration from one iteration is used to initialise the next, with $\mathcal{M}=\{\theta_{all},\theta_{1},\theta_{2},\theta_{3}\}$. Training the full model for additional epochs on the entire dataset degrades performance, with WER increasing from 9.48% to 10.28%, consistent with overfitting.
In contrast, iterative GRAPAM improves recognition: starting from speaker-embedding clustering with FFN partial fine-tuning, the best result is 9.28%, obtained at iteration 2 after applying LSM-based clustering with full fine-tuning. This gain is consistent with incremental exposure to complementary cluster structures across iterations, which may encourage more robust representations. After iteration 2, performance saturates and no further gains are observed, suggesting limited additional information in the available partitions and potential onset of overfitting. Overall, the results support iterative group-aware model merging as an effective mechanism for improving children’s ASR.
6 Conclusion and future work
In this work, we introduced GRAPAM, a parameter-efficient framework for adapting adult-pretrained ASR models to children’s speech via group-aware partial model merging. By integrating unsupervised clustering, partial fine-tuning, and parameter-level interpolation, GRAPAM reduces WER on MyST from 9.95% to 9.31%, achieving comparable or improved accuracy without multi-task training or additional data.
Several directions remain for future work. These include replacing uniform interpolation with adaptive weighting, learning merge coefficients on held-out validation data. Exploring stronger merging strategies, including clustering based on self-supervised speech representations and extending GRAPAM to other tasks such as pathological speech.
7 Acknowledgements
This work has been submitted to Interspeech 2026.
Work supported by Portuguese national funds through Fundação para a Ciência e a Tecnologia (FCT), with references UIDB/50021/2020 and 2022/12328/BD, as well as by the Portuguese Recovery and Resilience Plan (RRP) through project C644865762-00000008 (Accelerat.AI).